1.

A E 3 4 5 6 7 8 9 B D Budget Assumptions For the months of January, February, and March January February Selling price per unApplying Excel 14: Budgeted units to produce Which of the following formulas would be correct in cell D28?Multiple Choice (0 =+D22*D23+D24+D25-D27 =+$E$7*($F$10+$D$7)-F12 =+E7*(F10+D7)-F12 =(+E7*F10)+(D7-F12)2.

А F 3 4 5 6 7 8 9 10 11 B с Budget Assumptions For the months of January, February, and March January February Selling priceApplying Excel 14: Budgeted direct labor cost Which of the following formulas in cell D26 would yield the correct answer whenMultiple Choice =+$D$21*D22*D25 =+E21*D22*D25 (0) =+D21*D22*D25 =+$D$23*D25
Answer #1

Budgeted unit to produce = Closing stock + Current months budgeted sales – Opening stock

                                               = (next month budgeted sales*ratio of inventory to next month’s sales) +                                                                                                                                    Current months budgeted sales – Opening stock

                                               = (E7*F10)+(D7-F12)

                                               = (15,000*20%)+(12,000-2,400)

                                               = 12,600

Hence the answer is (E7*F10)+(D7-F12)

Budgeted direct labor cost = Budgeted unit to produce*Direct labor requirement per units*direct labor rate

                                                 = D21*D22*D25

                                                 = 12,600*2*18

                                                 = $ 453,600

Since the data of different months contain in different cell, there is no need of any absolute cell reference. Hence the answer is D21*D22*D25